Find the midpoint of the line segment with endpoints (-5,-2) and (7,6).
Answer by stanbon(75887)

Find the midpoint of the line segment with endpoints (-5,-2) and (7,6).
------
Let (x,y) be the midpoint.
--------
Equations:
x = (-5+7)/2 = 1
----
y = (6-2)/2 = 2
------------------
Cheers,
Stan H.
===============